nickpick1615:

PLZ HELP ME I NEED IT ON THIS QUESTION One link in a chain was made from a cylinder that has a radius of 3 cm and a height of 25 cm. How much plastic coating would be needed to coat the surface of the chain link? Use 3.14 for π. a chain link made from a cylinder that has been bent so that its ends are connected, creating a continuous loop 251.2 cm2 314 cm2 345.4 cm2 471 cm2

nickpick1615:

Id appreciate any help

nickpick1615:

can you help me?

AquarytheALPHA:

i can try

nickpick1615:

thank you

AquarytheALPHA:

Im trying to figure it out.

nickpick1615:

hello owen, and ok thank you and sorry for rushing

surjithayer:

surface area \[=2\pi r h=2 \times3.14\times 3 \times 25=3.14\times 225= 471 ~cm^2\]